Two angles of a triangle have the measure of 55 and 65. Which of the following could not be a measure of an exterior angle of the triangle?
Answer by Alan3354(69443)

Two angles of a triangle have the measure of 55 and 65. Which of the following could not be a measure of an exterior angle of the triangle?
---------
The 3 exterior angles are
180-55 = 125,
180-60 = 120 &
180-65 = 115